# The Ginsburg–Sands theorem and computability theory

### v1

@@ 6.2. $\mathsf{GST}_{1}$ and $\mathsf{RT}^{2}_{2}$
Liu [28, Theorem 1.5 and Corollary 1.6] established that $\mathsf{COH}$ admits PA avoidance, and $\mathsf{RT}^{1}_{2}$ admits strong PA avoidance. This means that for any set $X\not\gg\emptyset$̸ ∅, every $X$-computable instance of $\mathsf{COH}$ has a solution $Y$ such that $X\oplus Y\not\gg\emptyset$̸ ∅, and every instance of $\mathsf{RT}^{1}_{2}$ (computable from $X$ or not) has a solution $H$ such that $X\oplus H\not\gg\emptyset$̸ ∅. We can combine these facts to produce an $\omega$-model of $\Sigma^{0}_{2}\text{-}\mathsf{Subset}+\mathsf{COH}$, and hence (by Propositions 6.4 and 6.8) of $\mathsf{GST}_{1}$, that contains no set of PA degree, and therefore cannot satisfy $\mathsf{WKL}_{0}$. For completeness, we give the details. We define a sequence $X_{0}\leq_{\text{T}}X_{1}\leq_{\text{T}}\cdots$ of subsets of $\omega$, and define (the second-order part of) an $\omega$-model $\mathcal{M}$ to be $\{Y:(\exists s)(Y\leq_{\text{T}}X_{s})\}$. Let $X_{0}=\emptyset$, and suppose inductively that we have defined $X_{s}\not\gg\emptyset$̸ ∅ for some $s\in\omega$. First, suppose $s=2\langle e,t\rangle$, so that $e,t\leq s$ by standard conventions. If $\Phi_{e}^{X_{t}}$ defines a family $\vec{R}$ of subsets of $\omega$, then also $\vec{R}\leq_{\text{T}}X_{s}$, and we appeal to PA avoidance of $\mathsf{COH}$ to fix an infinite $\vec{R}$-cohesive set $Y$ such that $X_{s}\oplus Y\not\gg\emptyset$̸ ∅. We then let $X_{s+1}=X_{s}\oplus Y$. If $\Phi_{e}^{X_{t}}$ does not define a family, we simply let $X_{s+1}=X_{s}$. Now, suppose $s=2\langle e,t\rangle+1$. We consider $W_{e}^{X_{t}^{\prime}}$, the $e$th $\Sigma^{0}_{2}$ set relative to $X_{t}$, and regard it as the instance $c:\omega\to 2$ of $\mathsf{RT}^{1}_{2}$ defined by $c(x)=1$ if and only if $x\in W_{e}^{X_{t}^{\prime}}$. We appeal to strong PA avoidance of $\mathsf{RT}^{1}_{2}$ to fix an infinite homogeneous set $H$ for $c$ such that $X_{s}\oplus H\not\gg\emptyset$̸ ∅, noting that $H$ is an infinite subset of either $W_{e}^{X_{t}^{\prime}}$ or its complement. We then let $X_{s+1}=X_{s}\oplus H$. This completes the construction of $\mathcal{M}$. Since $X_{s}\not\gg\emptyset$̸ ∅ for all $s$, ⊧̸$\mathcal{M}\not\models\mathsf{WKL}_{0}$. To verify that $\mathcal{M}$ satisfies $\mathsf{COH}$, consider any instance $\vec{R}$ of $\mathsf{COH}$ in $\mathcal{M}$. Then $\vec{R}=\Phi_{e}^{X_{t}}$ for some $e$ and $t$, and a solution to $\vec{R}$ is thus computable from $X_{2\langle e,t\rangle}$ and so belongs to $\mathcal{M}$. Next, consider any instance of $\Sigma^{0}_{2}\text{-}\mathsf{Subset}$ in $\mathcal{M}$, meaning a set of the form $W_{e}^{X_{t}^{\prime}}$ for some $e$ and $t$. Then a solution to this instance is computable from $X_{2\langle e,t\rangle+1}$. ∎
@@
In particular, it follows that $\mathsf{GST}_{1}$ also admits cone avoidance, and so does not imply $\mathsf{ACA}_{0}$. For a direct proof of this result, which provides further insights into the computability-theoretic aspects of $T_{1}$ CSC spaces, see Benham [1].
@@ 6.3. Arithmetical bounds
Write $\mathcal{U}=\langle U_{n}:n\in\omega\rangle$. The proof breaks into three scenarios for finding the required subspace. First, consider the $\Sigma^{0}_{2}$ set $S=\{x\in X:(\exists n)[x\in U_{n}\wedge U_{n}\text{ is finite}]\}$. By Proposition 3.7, $x\in S$ if and only if $\{x\}\in\mathcal{U}$, so $S$ is a discrete subspace of $X$. If $S$ is infinite, we can fix an infinite $\Delta^{0}_{2}$ subset $Y$ of $S$, and this then forms an infinite $\Delta^{0}_{2}$ discrete subspace of $X$. So suppose $S$ is finite. By replacing $X$ with $X\smallsetminus S$, and each $U\in\mathcal{U}$ with $U\smallsetminus S$, we may assume without loss of generality that every nonempty element of $\mathcal{U}$ is infinite. (Since $S$ is finite, this change actually results in a computable subspace of $X$.) Now, suppose there exists a nonempty $U\in\mathcal{U}$ such that $U$ contains no $T_{2}$ pairs. Then $U$ forms an infinite computable pure $T_{1}$ subspace of $X$, so by Lemma 6.2 (3) it has an infinite computable subspace $Y$ with the cofinite topology, and this is of course also a subspace of $X$. For the remaining case, then, we can assume every nonempty $U\in\mathcal{U}$ contains a $T_{2}$ pair. We use $\emptyset^{\prime}$ to construct an infinite $\Delta^{0}_{2}$ discrete subspace of $X$. The key point we will use is that, given $m,n\in\omega$, asking whether $U_{n}=\emptyset$, whether $U_{n}\cap U_{m}=\emptyset$, or whether $U_{n}\subseteq U_{m}$ can all be answered by $\emptyset^{\prime}$, uniformly in $n$ and $m$. This has the following consequences. First, given $n$ such that $U_{n}\neq\emptyset$, $\emptyset^{\prime}$ can uniformly find a $T_{2}$ pair $x,y\in U_{n}$. And second, suppose we are given $i$ and $x,y\in U_{i}$ forming a $T_{2}$ pair. Then $\emptyset^{\prime}$ can uniformly find indices $m$ and $n$ such that $x\in U_{m}$, $y\in U_{n}$, $U_{m}\cap U_{n}=\emptyset$, and $U_{m},U_{n}\subseteq U_{i}$. To see this note, $\emptyset^{\prime}$ first uniformly finds $s$ and $t$ such that $x\in U_{s}$, $y\in U_{t}$, $U_{s}\cap U_{t}=\emptyset$. Then $m=k(x,i,s)$ and $n=k(x,i,t)$ have the desired properties. Indeed, we have $x\in U_{n}\subseteq U_{i}\cap U_{s}$ and $y\in U_{m}\subseteq U_{i}\cap U_{t}$, and $U_{n}\cap U_{m}\subseteq U_{s}\cap U_{t}=\emptyset$, as needed. We now construct $\Delta^{0}_{2}$ sequences of points $x_{0}<x_{1}<\cdots$ and $y_{0}<y_{1}<\cdots$ in $X$, and sequences of indices $m_{0},m_{1},\ldots$ and $n_{0},n_{1},\ldots$ such that, for all $i$:
  • $x_{i}\in X_{m_{i}}$;
@@
In particular, $x_{i}\in U_{m_{j}}$ if and only if $i=j$. Hence, if we set $Y=\{x_{i}:i\in\omega\}$, then $Y$ forms the desired $\Delta^{0}_{2}$ discrete subspace of $X$. To define our sequences, we proceed by recursion. Let $x_{0},y_{0}$ be an arbitrary $T_{2}$ pair in $X$, and let $m_{0}$ and $n_{0}$ be such that $x_{0}\in U_{m_{0}}$, $y_{0}\in U_{n_{0}}$, and $U_{m_{0}}\cap U_{n_{0}}=\emptyset$. Now fix $i$, and suppose $x_{i}$, $y_{i}$, $m_{i}$, and $n_{i}$ have been defined and satisfy the above properties. Since $U_{n_{i}}$ is nonempty, $\emptyset^{\prime}$ can uniformly find a $T_{2}$ pair $x,y\in U_{n_{i}}$. It can then uniformly find indices $m_{i+1}$ and $n_{i+1}$ such that $x\in U_{m_{i+1}}$, $y\in U_{n_{i+1}}$, $U_{m_{i+1}}\cap U_{n_{i+1}}=\emptyset$, and $U_{m_{i+1}},U_{n_{i+1}}\subseteq U_{n_{i}}$. Now $U_{m_{i+1}}$ and $U_{n_{i+1}}$ are nonempty, hence infinite, so we can fix the least $x_{i+1}>x_{i}$ in $U_{m_{i+1}}$, and the least $y_{i+1}>y_{i}$ in $U_{n_{i+1}}$. Clearly, $x_{i+1}$, $y_{i+1}$, $m_{i+1}$, and $n_{i+1}$ maintain the above properties. This completes the construction, and finishes the proof. ∎
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 7. Separating $\mathsf{GST}_{1}$ from $\mathsf{RT}^{2}_{2}$
The property we design here is new, and rather different in flavor from prior examples. It also has the unusual feature that it is $\Sigma^{1}_{1}$; prior examples have all been arithmetical.
Our argument divides into the following three lemmas. The first expresses a so-called “fairness property”, while the second and third express that this property is preserved under $\mathsf{D}^{2}_{2}$ and $\mathsf{COH}$. We recall that if $S,U\subseteq\omega$ are sets, then $S$ is $U$-immune if $U$ computes no infinite subset of $S$; $S$ is $U$-hyperimmune if $U$ computes no infinite set $\{x_{0}<x_{1}<\cdots\}\subseteq\omega$ such that $[x_{n},x_{n+1})\cap S\neq\emptyset$ for all $n$. Being $U$-hyperimmune implies being $U$-immune. If $S$ and $\overline{S}$ are both $U$-immune, or both $U$-hyperimmune, then $S$ is $U$-bi-immune, respectively, $U$-bi-hyperimmune.
@@
We construct an $\omega$-model by applying Lemma 7.2 and then iterating and dovetailing Lemmas 7.3 and 7.4. To this end, we define a chain of sets $X_{0}\leq_{\text{T}}X_{1}\leq_{\text{T}}\cdots$ and let our model be =$\mathcal{M}=\{Y:(\exists s)(Y\leq_{\text{T}}X_{s})\}$, as usual. We proceed by induction. Let $X_{0}=\emptyset$, and let $S$, $L_{0}$, and $P_{0}$ be as given by Lemma 7.2. Notice that, if we take $X=X_{0}$, $L=L_{0}$, and $P=P_{0}$, then $X$, $S$, $L$, and $P$ satisfy the hypotheses of Lemmas 7.3 and 7.4. Next, fix $s\in\omega$, and suppose that for each $t\leq s$ we have defined $X_{t}$, $L_{t}$, and $P_{t}$ such that, if we take $X=X_{t}$, $L=L_{t}$, and $P=P_{t}$, then $X$, $S$, $L$, and $P$ satisfy the hypotheses of Lemmas 7.3 and 7.4. Say $s$ is $2\langle e,t\rangle$ or $12\langle e,t\rangle+$ 1, so that $e,t\leq s$ by standard conventions. In the first case, we check if $\Phi_{e}^{X_{t}^{\prime}}$ defines a set $A$, and in the second, we check if $\Phi_{e}^{X_{t}}$ defines a family of sets $\vec{R}$. If not, we let $X_{s+1}=X_{s}$, $L_{s+1}=L_{s}$, and $P_{s+1}=P_{s}$ in either case. Otherwise, we have $A\leq_{\text{T}}X_{s}^{\prime}$ or $\vec{R}\leq_{\text{T}}X_{s}$, respectively. We can thus apply Lemma 7.3 in the first case, or Lemma 7.4 in the second, to obtain sets $G$, $L_{G}$, and $P_{G}$. (So, in the first case, we have that $G$ is an infinite subset of $A$ or $\overline{A}$, and in the second, that $G$ is an infinite $\vec{R}$-cohesive set.) In either case, as noted above, we have that $X_{s}\oplus G$ does not compute any infinite subset of $S$ or $\overline{S}$. We now set $X_{s+1}=X_{s}\oplus G$, $L_{s+1}=L_{G}$ and $P_{s+1}=P_{G}$, and notice that, if we take $X=X_{s+1}$, $L=L_{s+1}$, and $P=P_{s+1}$, then we again satisfy the hypotheses of Lemmas 7.3 and 7.4. So the inductive conditions are maintained. Let $\mathcal{M}$ be the model specified above. By construction, if $A$ is any instance of $\mathsf{D}^{2}_{2}$ in $\mathcal{M}$ then $A=\Phi_{e}^{X_{t}^{\prime}}$ for $e$ and $t$, and then some solution $G$ to $A$ is computable from $X_{2\langle e,t\rangle}$ and so belongs to $\mathcal{M}$. Similarly, if $\vec{R}$ is any instance of $\mathsf{COH}$ in $\mathcal{M}$ then $\vec{R}=\Phi_{e}^{X_{t}}$ for some for $e$ and $t$, and now a solution $G$ to $\vec{R}$ is computable $X_{2\langle e,t\rangle+1}$. Hence, $\mathcal{M}$ satisfies $\mathsf{D}^{2}_{2}+\mathsf{COH}$, and therefore $\mathsf{RT}^{2}_{2}$. On the other hand, by induction, no $X_{s}$ computes any infinite subset of $S$ or $\overline{S}$, so $S$ is a computable instance of $\Sigma^{0}_{2}\text{-}\mathsf{Subset}$ with no solution in $\mathcal{M}$. Hence, $\mathcal{M}$ does not satisfy $\Sigma^{0}_{2}\text{-}\mathsf{Subset}$. ∎
@@
Proof of Lemma 7.2.
We enumerate $S$ by stages, using a $\emptyset^{\prime}$ oracle. For $L_{0}$, we can take any low set of PA degree, but the set $P_{0}$ will have to be constructed alongside $S$. For ease of notation, let $T=T^{2}_{\emptyset^{\prime}}$. We will approximate an element of this tree during the construction, and take this to be $P_{0}$ at the end. We aim to satisfy the following requirements for all $e\in\omega$:
$\mathcal{L}_{e}$ ::: $\Phi^{L_{0}}_{e}$ is not the characteristic function of an infinite set or

### v2

@@ 6.2. $\mathsf{GST}_{1}$ and $\mathsf{RT}^{2}_{2}$
Liu [30, Theorem 1.5 and Corollary 1.6] established that $\mathsf{COH}$ admits PA avoidance, and $\mathsf{RT}^{1}_{2}$ admits strong PA avoidance. This means that for any set $X\not\gg\emptyset$̸ ∅ (i.e., for any set $X$ not of PA degree), every $X$-computable instance of $\mathsf{COH}$ has a solution $Y$ such that $X\oplus Y\not\gg\emptyset$̸ ∅, and every instance of $\mathsf{RT}^{1}_{2}$ (computable from $X$ or not) has a solution $H$ such that $X\oplus H\not\gg\emptyset$̸ ∅. We can combine these facts to produce an $\omega$-model of $\Sigma^{0}_{2}\text{-}\mathsf{Subset}+\mathsf{COH}$, and hence (by Propositions 6.4 and 6.8) of $\mathsf{GST}_{1}$, that contains no set of PA degree, and therefore cannot satisfy $\mathsf{WKL}_{0}$. For completeness, we give the details. We define a sequence $X_{0}\leq_{\text{T}}X_{1}\leq_{\text{T}}\cdots$ of subsets of $\omega$, and define (the second-order part of) an $\omega$-model $\mathcal{M}$ to be $\{Y:(\exists s)(Y\leq_{\text{T}}X_{s})\}$. Let $X_{0}=\emptyset$, and suppose inductively that we have defined $X_{s}\not\gg\emptyset$̸ ∅ for some $s\in\omega$. First, suppose $s=2\langle e,t\rangle$, so that $e,t\leq s$ by standard conventions. If $\Phi_{e}^{X_{t}}$ defines a family $\vec{R}$ of subsets of $\omega$, then also $\vec{R}\leq_{\text{T}}X_{s}$, and we appeal to PA avoidance of $\mathsf{COH}$ to fix an infinite $\vec{R}$-cohesive set $Y$ such that $X_{s}\oplus Y\not\gg\emptyset$̸ ∅. We then let $X_{s+1}=X_{s}\oplus Y$. If $\Phi_{e}^{X_{t}}$ does not define a family, we simply let $X_{s+1}=X_{s}$. Now, suppose $s=2\langle e,t\rangle+1$. We consider $W_{e}^{X_{t}^{\prime}}$, the $e$th $\Sigma^{0}_{2}$ set relative to $X_{t}$, and regard it as the instance $c:\omega\to 2$ of $\mathsf{RT}^{1}_{2}$ defined by $c(x)=1$ if and only if $x\in W_{e}^{X_{t}^{\prime}}$. We appeal to strong PA avoidance of $\mathsf{RT}^{1}_{2}$ to fix an infinite homogeneous set $H$ for $c$ such that $X_{s}\oplus H\not\gg\emptyset$̸ ∅, noting that $H$ is an infinite subset of either $W_{e}^{X_{t}^{\prime}}$ or its complement. We then let $X_{s+1}=X_{s}\oplus H$. This completes the construction of $\mathcal{M}$. Since $X_{s}\not\gg\emptyset$̸ ∅ for all $s$, ⊧̸$\mathcal{M}\not\models\mathsf{WKL}_{0}$. To verify that $\mathcal{M}$ satisfies $\mathsf{COH}$, consider any instance $\vec{R}$ of $\mathsf{COH}$ in $\mathcal{M}$. Then $\vec{R}=\Phi_{e}^{X_{t}}$ for some $e$ and $t$, and a solution to $\vec{R}$ is thus computable from $X_{2\langle e,t\rangle}$ and so belongs to $\mathcal{M}$. Next, consider any instance of $\Sigma^{0}_{2}\text{-}\mathsf{Subset}$ in $\mathcal{M}$, meaning a set of the form $W_{e}^{X_{t}^{\prime}}$ for some $e$ and $t$. Then a solution to this instance is computable from $X_{2\langle e,t\rangle+1}$. ∎
@@
In particular, it follows that $\mathsf{GST}_{1}$ also does not imply $\mathsf{ACA}_{0}$. In fact, by using strong cone avoidance of $\mathsf{RT}^{1}_{2}$ in place of strong PA avoidance in the proof above, we can deduce that $\mathsf{GST}_{1}$ admits cone avoidance. For another proof of this result, which provides further insights into the computability-theoretic aspects of $T_{1}$ CSC spaces, see Benham [1].
@@ 6.3. Arithmetical bounds
Write $\mathcal{U}=\langle U_{n}:n\in\omega\rangle$. The proof breaks into three scenarios for finding the required subspace. First, consider the $\Sigma^{0}_{2}$ set $S=\{x\in X:(\exists n)[x\in U_{n}\wedge U_{n}\text{ is finite}]\}$. By Proposition 3.7, $x\in S$ if and only if $\{x\}\in\mathcal{U}$, so $S$ is a discrete subspace of $X$. If $S$ is infinite, we can fix an infinite $\Delta^{0}_{2}$ subset $Y$ of $S$, and this then forms an infinite $\Delta^{0}_{2}$ discrete subspace of $X$.
So suppose $S$ is finite. By replacing $X$ with $X\smallsetminus S$, and each $U\in\mathcal{U}$ with $U\smallsetminus S$, we may assume without loss of generality that every nonempty element of $\mathcal{U}$ is infinite. (Since $S$ is finite, this change actually results in a computable subspace of $X$.) Now, suppose there exists a nonempty $U\in\mathcal{U}$ such that $U$ contains no $T_{2}$ pairs. Then $U$ forms an infinite computable pure $T_{1}$ subspace of $X$, so by Lemma 6.2 (3) it has an infinite computable subspace $Y$ with the cofinite topology, and this is of course also a subspace of $X$.
For the remaining case, then, we can assume every nonempty $U\in\mathcal{U}$ contains a $T_{2}$ pair. We use $\emptyset^{\prime}$ to construct an infinite $\Delta^{0}_{2}$ discrete subspace of $X$. The key point we will use is that, given $m,n\in\omega$, asking whether $U_{n}=\emptyset$, whether $U_{n}\cap U_{m}=\emptyset$, or whether $U_{n}\subseteq U_{m}$ can all be answered by $\emptyset^{\prime}$, uniformly in $n$ and $m$. This has the following consequences. First, given $n$ such that $U_{n}\neq\emptyset$, $\emptyset^{\prime}$ can uniformly find a $T_{2}$ pair $x,y\in U_{n}$. And second, suppose we are given $i$ and $x,y\in U_{i}$ forming a $T_{2}$ pair. Then $\emptyset^{\prime}$ can uniformly find indices $m$ and $n$ such that $x\in U_{m}$, $y\in U_{n}$, $U_{m}\cap U_{n}=\emptyset$, and $U_{m},U_{n}\subseteq U_{i}$. To see this note, $\emptyset^{\prime}$ first uniformly finds $s$ and $t$ such that $x\in U_{s}$, $y\in U_{t}$, $U_{s}\cap U_{t}=\emptyset$. Then $m=k(x,i,s)$ and $n=k(x,i,t)$ have the desired properties. Indeed, we have $x\in U_{n}\subseteq U_{i}\cap U_{s}$ and $y\in U_{m}\subseteq U_{i}\cap U_{t}$, and $U_{n}\cap U_{m}\subseteq U_{s}\cap U_{t}=\emptyset$, as needed.
We now construct $\Delta^{0}_{2}$ sequences of points $x_{0}<x_{1}<\cdots$ and $y_{0}<y_{1}<\cdots$ in $X$, and sequences of indices $m_{0},m_{1},\ldots$ and $n_{0},n_{1},\ldots$ such that, for all $i$:
  • $x_{i}\in U_{m_{i}}$;
@@
In particular, $x_{i}\in U_{m_{j}}$ if and only if $i=j$. Hence, if we set $Y=\{x_{i}:i\in\omega\}$, then $Y$ forms the desired $\Delta^{0}_{2}$ discrete subspace of $X$.
To define our sequences, we proceed by recursion. Let $x_{0},y_{0}$ be an arbitrary $T_{2}$ pair in $X$, and let $m_{0}$ and $n_{0}$ be such that $x_{0}\in U_{m_{0}}$, $y_{0}\in U_{n_{0}}$, and $U_{m_{0}}\cap U_{n_{0}}=\emptyset$. Now fix $i$, and suppose $x_{i}$, $y_{i}$, $m_{i}$, and $n_{i}$ have been defined and satisfy the above properties. Since $U_{n_{i}}$ is nonempty, $\emptyset^{\prime}$ can uniformly find a $T_{2}$ pair $x,y\in U_{n_{i}}$. It can then uniformly find indices $m_{i+1}$ and $n_{i+1}$ such that $x\in U_{m_{i+1}}$, $y\in U_{n_{i+1}}$, $U_{m_{i+1}}\cap U_{n_{i+1}}=\emptyset$, and $U_{m_{i+1}},U_{n_{i+1}}\subseteq U_{n_{i}}$. Now $U_{m_{i+1}}$ and $U_{n_{i+1}}$ are nonempty, hence infinite, so we can fix the least $x_{i+1}>x_{i}$ in $U_{m_{i+1}}$, and the least $y_{i+1}>y_{i}$ in $U_{n_{i+1}}$. Clearly, $x_{i+1}$, $y_{i+1}$, $m_{i+1}$, and $n_{i+1}$ maintain the above properties. This completes the construction, and finishes the proof. ∎
@@
In spite of the curious distinction between $\mathsf{GST}_{1}$ and $\mathsf{RT}^{2}_{2}$ highlighted by Theorem 6.12, there are other aspects in which $\mathsf{GST}_{1}$ and $\mathsf{RT}^{2}_{2}$ behave quite similarly. One of these is PA avoidance, which holds for $\mathsf{RT}^{2}_{2}$ by Liu’s theorem (discussed in the introduction) and for $\mathsf{GST}_{1}$ by Corollary 6.11 above. Another is the following. Cholak, Jockusch, and Slaman [2, Theorem 3.1] showed that every computable instance of $\mathsf{RT}^{2}_{2}$ has a solution $H$ which is low2 (i.e., satisfies $H^{\prime\prime}\leq_{\text{T}}\emptyset^{\prime\prime}$). In fact, they offered two proofs of this result, one by controlling the first jump of $H$ during its construction ([2], Section 4), the other by controlling its second jump ([2], Section 5). We can adapt the latter proof to obtain the analogous result for $\mathsf{GST}_{1}$.
@@
In [2, Section 5.2], it is shown that if $A$ is any $\Delta^{0}_{2}$ set then there exists an infinite subset of $A$ or $\overline{A}$ which is low2. The authors comment in the proof (Section 5.2.1 on p. 19) that, except for one moment, their argument also works for $A$ being $\Delta^{0}_{3}$. This moment is when they consider a certain sentence of the form
@@
Let $\langle X,\mathcal{U},k\rangle$ be an infinite computable $T_{1}$ CSC space. Then $X$ has an infinite low2 subspace which is discrete or has the cofinite topology.
@@
Fix $\langle X,\mathcal{U},k\rangle$. Jockusch and Stephan [27, Theorem 2.1] showed that every computable instance of $\mathsf{COH}$ has a low2 solution. So let $Y$ be an infinite low2 $\mathcal{U}$-cohesive set. Then $Y$ is stable as a subspace of $X$, and we let $Y$ be the $\Sigma^{0}_{2}(Y)$ set of all $x\in Y$ such that $(\exists U\in\mathcal{U})[U=\{x\}]$. By the proof of Proposition 6.8, every infinite subset of $S$ is a discrete subspace of $Y$ (and hence of $X$) and every infinite subset of $Y\smallsetminus S$ is a subspace with the cofinite topology. Apply Theorem 6.14 to find an infinite subset $Z$ of $S$ or $Y\smallsetminus S$ which is low2 over $Y$. Then $Z$ is itself low2, which completes the proof. ∎
@@
There exists an $\omega$-model of $\mathsf{GST}_{1}$ consisting entirely of low2 sets.
@@ 7. Separating $\mathsf{GST}_{1}$ from $\mathsf{RT}^{2}_{2}$
Our argument divides into the following three lemmas. The first expresses a so-called “fairness property”, while the second and third express that this property is preserved under $\mathsf{D}^{2}_{2}$ and $\mathsf{COH}$. We recall that if $S,U\subseteq\omega$ are sets, then $S$ is $U$-immune if $U$ computes no infinite sequence of $S$; $S$ is $U$-hyperimmune if $U$ computes no infinite sequence $x_{0}<x_{1}<\cdots$ of numbers such that $[x_{n},x_{n+1})\cap S\neq\emptyset$ for all $n$. Say $\Phi^{U}_{e}$ is a non-hyperimmunity witness for $S$ if it is the characteristic function of such a sequence. Thus, a set $S$ is $U$-hyperimmune if and only if $\Phi^{U}_{e}$ is not a non-hyperimmunity witness for $S$, for all $e$. Being $U$-hyperimmune implies being $U$-immune. If $S$ and $\overline{S}$ are both $U$-immune, or both $U$-hyperimmune, then $S$ is $U$-bi-immune, respectively, $U$-bi-hyperimmune. If $S$ computes no $U$-hyperimmune set then it is said to be of hyperimmune-free degree relative to $U$, and this is equivalent to every $S$-computable function being dominated by a $U$-computable function. (See, e.g., [11, Section 2.17] for details on these notions.)
@@
We construct an $\omega$-model by applying Lemma 7.2 and then iterating and dovetailing Lemmas 7.3 and 7.4. To this end, we define a chain of sets $X_{0}\leq_{\text{T}}X_{1}\leq_{\text{T}}\cdots$ and let our model be =$\mathcal{M}=\{Y:(\exists s)(Y\leq_{\text{T}}X_{s})\}$, as usual. We proceed by induction.
Let $X_{0}=\emptyset$, and let $S$, $L_{0}$, and $P_{0}$ be as given by Lemma 7.2. Notice that, if we take $X=X_{0}$, $L=L_{0}$, and $P=P_{0}$, then $X$, $S$, $L$, and $P$ satisfy the hypotheses of Lemmas 7.3 and 7.4.
Next, fix $s\in\omega$, and suppose that for each $t\leq s$ we have defined $X_{t}$, $L_{t}$, and $P_{t}$ such that, if we take $X=X_{t}$, $L=L_{t}$, and $P=P_{t}$, then $X$, $S$, $L$, and $P$ satisfy the hypotheses of Lemmas 7.3 and 7.4. Say $s$ is $2\langle e,t\rangle$ or $12\langle e,t\rangle+$ 1, so that $e,t\leq s$ by standard conventions. In the first case, we check if $\Phi_{e}^{X_{t}^{\prime}}$ defines a set $A$, and in the second, we check if $\Phi_{e}^{X_{t}}$ defines a family of sets $\vec{R}$. If not, we let $X_{s+1}=X_{s}$, $L_{s+1}=L_{s}$, and $P_{s+1}=P_{s}$ in either case. Otherwise, we have $A\leq_{\text{T}}X_{s}^{\prime}$ or $\vec{R}\leq_{\text{T}}X_{s}$, respectively. We can thus apply Lemma 7.3 in the first case, or Lemma 7.4 in the second, to obtain sets $G$, $L_{G}$, and $P_{G}$. (So, in the first case, we have that $G$ is an infinite subset of $A$ or $\overline{A}$, and in the second, that $G$ is an infinite $\vec{R}$-cohesive set.) In either case, as noted above, we have that $X_{s}\oplus G$ does not compute any infinite subset of $S$ or $\overline{S}$. We now set $X_{s+1}=X_{s}\oplus G$, $L_{s+1}=L_{G}$ and $P_{s+1}=P_{G}$, and notice that, if we take $X=X_{s+1}$, $L=L_{s+1}$, and $P=P_{s+1}$, then we again satisfy the hypotheses of Lemmas 7.3 and 7.4. So the inductive conditions are maintained.
Let $\mathcal{M}$ be the model specified above. By construction, if $A$ is any instance of $\mathsf{D}^{2}_{2}$ in $\mathcal{M}$ then $A=\Phi_{e}^{X_{t}^{\prime}}$ for $e$ and $t$, and then some solution $G$ to $A$ is computable from $X_{2\langle e,t\rangle}$ and so belongs to $\mathcal{M}$. Similarly, if $\vec{R}$ is any instance of $\mathsf{COH}$ in $\mathcal{M}$ then $\vec{R}=\Phi_{e}^{X_{t}}$ for some for $e$ and $t$, and now a solution $G$ to $\vec{R}$ is computable $X_{2\langle e,t\rangle+1}$. Hence, $\mathcal{M}$ satisfies $\mathsf{D}^{2}_{2}+\mathsf{COH}$, and therefore $\mathsf{RT}^{2}_{2}$. On the other hand, by induction, no $X_{s}$ computes any infinite subset of $S$ or $\overline{S}$, so $S$ is a computable instance of $\Sigma^{0}_{2}\text{-}\mathsf{Subset}$ with no solution in $\mathcal{M}$. Hence, $\mathcal{M}$ does not satisfy $\Sigma^{0}_{2}\text{-}\mathsf{Subset}$. ∎
@@
For the remainder of this section, we follow the convention that if $\Phi^{F}(x)\operatorname{\downarrow}$ for some Turing functional $\Phi$, finite set $F$, and $x\in\omega$, then the use of this computation is bounded by $\max F$. For a set $U$, let $T^{2}_{U}\subseteq 2^{<\omega}$ be the standard $U$-computable tree whose paths are precisely the DNC2 functions relative to $U$. By well-known results of Jockusch and Soare [26] and Solovay (unpublished), $X\gg U$ if and only if $X\equiv_{\text{T}}f$ for some $f\in[T^{2}_{U}]$; see [11, Theorem 2.22.2].
Our original proof of Lemma 7.2 built $S$, $L_{0}$, and $P_{0}$ directly, via a slightly more complicated construction than the one that follows. We thank the anonymous referee for suggesting the more streamlined version below.
Proof of Lemma 7.2.
We let $L_{0}$ be any low set of PA degree, and $P_{0}\gg L_{0}^{\prime}$ any set of hyperimmune-free degree relative to $L_{0}^{\prime}$. These exist, respectively, by applying the low basis theorem to $T^{2}_{\emptyset}$, and the hyperimmune-free basis theorem, relativized to $L_{0}^{\prime}$, to $T^{2}_{L_{0}^{\prime}}$ (Jockusch and Soare [26, Theorems 2.2 and 2.4]). We construct a $\Sigma^{0}_{2}$ set $S$ such that $S$ is $L_{0}$-hyperimmune and $\overline{S}$ is $L_{0}^{\prime}$-hyperimmune. Since $P_{0}$ is of hyperimmune-free degree relative to $L_{0}^{\prime}$, this implies that $\overline{S}$ is also $P_{0}$-hyperimmune, as needed. (To see this, fix any $P_{0}$-computable sequence $x_{0}<x_{1}<\cdots$. Since $P_{0}$ is of hyperimmune-free degree relative to $L_{0}^{\prime}$, there is a function $g\leq_{\text{T}}L_{0}^{\prime}$ such that $g(n)\geq x_{n}$ for all $n$. Define a function $h$ by $h(0)=x_{0}$ and $h(n+1)=g(h(n)+1)$ for all $n$. Then $h\leq_{\text{T}}g\leq_{\text{T}}L_{0}^{\prime}$, and we have $h(n+1)\geq x_{h(n)+1}>x_{h(n)}\geq h(n)$ for all $n$. Since $\overline{S}$ is $L_{0}^{\prime}$-hyperimmune, there exists an $n$ such that $[h(n),h(n+1))\subseteq S$, and hence also $[x_{h(n)},x_{h(n)+1})\subseteq S$.)
We enumerate $S$ by stages, using $L_{0}^{\prime}$ as an oracle. Thus $S$ will be $\Sigma^{0}_{1}(L_{0}^{\prime})$, and so $\Sigma^{0}_{2}$ since $L_{0}$ is low. We aim to satisfy the following requirements for all $e\in\omega$:
$\mathcal{R}_{e}$ ::: $\Phi^{L_{0}}_{e}$ is not the characteristic function of an infinite set or
@@
$\mathcal{Q}_{e}$ ::: $\Phi^{L_{0}^{\prime}}_{e}$ is not the characteristic function of an infinite set or